OpenStudy (anonymous):

write each decimal as a fraction or mixed number 6.9

OpenStudy (anonymous):

multiply it by 1 \[1=\frac{10}{10}\]

OpenStudy (anonymous):

so for that the correct answer would be 6=9over 10

OpenStudy (anonymous):

can you rewrite that, the = is confusing me

OpenStudy (anonymous):

no = sorry

OpenStudy (anonymous):

improper fraction form would be \[\frac{69}{10}\] any idea how to turn that into a mixed number?

OpenStudy (anonymous):

so if I am dividing the fraction for example 4 over 5 is it 0.9

OpenStudy (anonymous):

.69

OpenStudy (anonymous):

4/5 = .8 4/5=8/10=.8 9/10 = .9

OpenStudy (anonymous):

how did you get that answer please show

OpenStudy (anonymous):

whenever you're dividing by 10, you move the decimal place over to the left by 1 when ever you're multiplying by 10 you move the decimal place over to the right by 1 or if you know long division, you can think of it like that

OpenStudy (anonymous):

to determine mixed numbers, you think of the decimal seperately example 6.4= 6 .4(10/10) \[6.4= 6 \frac{4}{10}\]